# Demystify Protein Generation with Hierarchical Conditional Diffusion Models

### v1

@@ 3.2 Evaluation of Conditional Consistency
To validate the effectiveness of Protein-MMD and other metrics, we apply the evaluation metrics on the Enzyme Commission (EC) dataset, which categorizes proteins based on the reactions they catalyze using EC numbers. We focus on seven classes from the first EC number category for our conditional generation task. In Table 1, we compare three evaluation metrics: MMD (considering only sequence statistics as presented in (Kucera et al., 2022)), Protein-FID (using ESM2 in place of Inception for protein generation), and Protein-MMD. All metrics are used to compute the Accuracy, Mean Reciprocal Rank (MRR), and Normalized Mean Rank (NMR) scores to evaluate how it performs in evaluating the conditional consistency. As observed, Protein-MMD outperforms both MMD and Protein-FID across all evaluation metrics. The higher accuracy and MRR scores indicate that Protein-MMD better captures the conditional consistency of the proteins in the test set. The lower NMR score further demonstrates that Protein-MMD ranks the correct target class higher in comparison to other metrics, validating its effectiveness in conditional protein generation tasks. While Protein-MMD proves to be the most effective metric according to our framework, we acknowledge the widespread use of FID in generative modeling tasks. Therefore, we will continue to report Protein-FID results alongside Protein-MMD in subsequent experiments.
4 Experiments
4.1 Experimental Setup
We compared our model against several baselines, each representing distinct approaches to protein generation. ProteoGAN (Kucera et al., 2022) is a GAN-based method, while ESM2 (Lin et al., 2023) and ProstT5 (Heinzinger et al., 2023) are Transformer-based language models specifically designed for protein sequence modeling. ProteinMPNN (Dauparas et al., 2022) and LatentDiff (Fu et al., 2024), on the other hand, are graph-based models, with LatentDiff also incorporating a diffusion-based framework, specifically using a latent diffusion approach. For each model, we evaluate the performance using both diversity metrics (TM-score, RMSD, and Seq.ID) and conditional consistency metrics (Protein-MMD and Protein-FID). Higher RMSD, lower TM-score, and lower Seq.ID indicate higher diversity, while lower Protein-MMD and Protein-FID values signify higher conditional consistency between the generated and real protein distributions.
Our model leverages the esm2_t33_650M_UR50D model from ESM2 (Lin et al., 2023) as the amino acid-level encoder. To construct the Protein Variational Auto-encoder model, we set the latent dimension to 384, and the decoder is composed of 8 Transformer (Vaswani, 2017) decoder blocks, each equipped with an 8-head self-attention mechanism. The Protein Variational Auto-encoder model is trained with a learning rate of $10^{-4}$, using a combination of mean squared error (MSE) and cross-entropy as the loss functions. To regulate the latent vector distribution, we apply a KL divergence loss with a weight of $10^{-5}$. We experimented with 128, 256, and 512 as the maximum sequence lengths. For the diffusion model, we modify the DiT-B architecture from DiT (Peebles & Xie, 2023b), which consists of 12 DiT blocks and uses a hidden size of 768. The DiT model is trained from scratch with a learning rate of $10^{-4}$ and includes a weight decay of $10^{-5}$.
4.2 Results and Analysis
Figure 5: Consistency in the latent space.
@@ 4.2 Results and Analysis
Impact of Maximum Sequence Length. In previous studies on protein de novo design, existing works usually employ a maximum sequence length of 128 (Fu et al., 2024). However, through our experiments, we observed that for conditional generation tasks, shorter sequence lengths fail to fully leverage the conditional information, which in turn results in lower conditional consistency metrics. To address this, we constructed models with three different maximum sequence lengths: 128, 256, and 512, and investigated the impact of maximum length on the model’s ability to maintain conditional consistency.
@@
Case Study. To further demonstrate the superiority of our model on the GO dataset, particularly regarding conditional consistency, we conducted a fine-grained case study comparing our method with the best baseline ProteoGAN. While our model leads on the EC dataset, it ranks second to ProteoGAN on the GO dataset in terms of conditional consistency metrics. We utilized an in-silico evaluation to perform a fine-grained analysis of the generated protein sequences. By employing a trained ESM-MLP classifier on the GO dataset, we assessed each generated protein’s adherence to the specified GO terms using the Intersection over Union (IoU) (Rezatofighi et al., 2019; Gao et al., 2023b).
As shown in Table 4, our method exhibits a lower average IoUmean compared to ProteoGAN, aligning with earlier results in Table 2. However, it achieves a higher IoUmax, indicating a greater potential for generating high-quality samples that closely match the desired GO annotations. Figure 6 illustrates the distribution of IoU scores. While ProteoGAN’s samples are concentrated around medium quality, our method generates a broader range of samples, including those with higher IoU scores. This suggests that our model, despite a lower average performance, is more capable of producing proteins with superior conditional consistency.
@@ Appendix A Appendices
A.1 Related Works
De novo protein design methods are dedicated to identify novel proteins with the desired structure and function properties (Watson et al., 2023; Huang et al., 2016; Frey et al., 2024; Mao et al., 2024; Komorowska et al., 2024; Gao et al., 2024b). Recent advancements in machine learning have enabled a generative model to accelerate key steps in the discovery of novel molecular structures and drug design (Gao et al., 2023a; Lu et al., 2022; Pei et al., 2023; Liu et al., 2022). A prior step of generate models to the representation of proteins (Zhang et al., 2023; Zhou et al., 2023; Gong et al., 2024; Zhao et al., 2024; Liu et al., 2023; Somnath et al., 2021; Jamasb et al., 2024). The majority of representation learning for protein is to represent a protein as a sequence of amino acids (Chen et al., 2023; Moreta et al., 2022; Lee et al., 2024; Fan et al., 2023). Considering the spatial information is important to the property of a protein, many works resort to a graph model for a comprehensive presentation with the structure information (Ingraham et al., 2019; Huang et al., 2024a). In general, each node on the graph is an amino acid and the edge is decided by the distance between two nodes (Aykent & Xia, 2022; Xia & Ku, 2021; Hladis et al., 2023). Despite the power of graph models, the relation information in a 3-dimensional space captures the multi-level structure such as the angle between two edges. A line of research works explore the protein structure in 3D space (Hermosilla et al., 2021; Huang et al., 2024b; Wang et al., 2023; Zhong et al., 2020; Peng et al., 2022; Liu et al., 2022; Huang et al., 2024a). Recently, large language models (LLMs) have also been introduced to model the sequence (Meier et al., 2021; Lin et al., 2023; Hu et al., 2022; Su et al., 2024) inspired by the success of natural language processing.
Capitalizing on the power of generative models such as Generative Adversarial Networks (GANs) and diffusion models, deep generative modeling has shown its potential for fast generation of new and viable protein structures. (Anand & Huang, 2018) has applied GANs to the task of generating protein structures by encoding protein structures in terms of pairwise distances on the protein backbone. Diffusion models (Song et al., 2021; Jiang et al., 2023; Gao et al., 2024a) have emerged as a powerful tool for graph-structured diffusion processes (Klarner et al., 2024). FrameDiff has been proposed for monomer backbone generation and it can generate designable monomers up to 500 amino acids (Yim et al., 2023). NOS is another diffusion model that generates protein sequences with high likelihood by taking many alternating steps in the continuous latent space of the model (Gruver et al., 2023).

### v2

@@ 3.2 Results and Analysis
Figure 5: Consistency.
Impact of Maximum Sequence Length. In previous studies on protein de novo design, existing works usually employ a maximum sequence length of 128 [fu2024latent]. However, through our experiments, we observed that for conditional generation tasks, shorter sequence lengths fail to fully leverage the conditional information, which in turn results in lower conditional consistency metrics. To address this, we constructed models with three different maximum sequence lengths: 128, 256, and 512, and investigated the impact of maximum length on the model’s ability to maintain conditional consistency.
@@
Case Study. To further demonstrate the superiority of our model on the GO dataset, particularly regarding conditional consistency, we conducted a fine-grained case study comparing our method with the best baseline ProteoGAN. We utilized an in-silico evaluation to perform a fine-grained analysis of the generated protein sequences. By employing a trained ESM-MLP classifier on the GO dataset, we assessed each generated protein’s adherence to the specified GO terms using the Intersection over Union (IoU) [DBLP:conf/cvpr/RezatofighiTGS019]. As shown in Table 4, our method exhibits a lower average IoUmean compared to ProteoGAN, aligning with earlier results in Table 2. However, it achieves a higher IoUmax, indicating a greater potential for generating high-quality samples that closely match the desired GO annotations. Figure 6 illustrates the distribution of IoU scores. While ProteoGAN’s samples are concentrated around medium quality, our method generates a broader range of samples, including those with higher IoU scores. This suggests that our model, despite a lower average performance, is more capable of producing proteins with superior conditional consistency.
4 Related Works
De novo protein design methods are dedicated to identify novel proteins with the desired structure and function properties [watson2023novo, huang2016coming, DBLP:conf/iclr/FreyBZKLHWRBCLG24, DBLP:conf/iclr/MaoZSSW0S24, DBLP:conf/iclr/KomorowskaMDVLJ24, DBLP:conf/iclr/GaoSLL0024]. Recent advancements in machine learning have enabled a generative model to accelerate key steps in the discovery of novel molecular structures and drug design [DBLP:conf/nips/GaoQTJRLLML23, DBLP:conf/nips/LuWZRLZ22, DBLP:conf/nips/PeiGWZX [^0] L023, DBLP:conf/icml/LiuLUMJ22]. A prior step of generate models to the representation of proteins [DBLP:conf/iclr/ZhangXJCLD023, DBLP:conf/iclr/ZhouFZBY23, DBLP:conf/icml/GongKLC0D24]. The majority of representation learning for protein is to represent a protein as a sequence of amino acids [DBLP:conf/iclr/ChenGDCW0WEDK23, DBLP:conf/iclr/MoretaRAHTH22, DBLP:conf/iclr/LeeY0K24, DBLP:conf/iclr/FanW0K23]. Considering the spatial information is important to the property of a protein, many works resort to a graph model for a comprehensive presentation with the structure information [DBLP:conf/iclr/IngrahamGBJ19, DBLP:conf/aaai/HuangLWSLZLGZL24]. In general, each node on the graph is an amino acid and the edge is decided by the distance between two nodes [DBLP:conf/kdd/AykentX22, DBLP:conf/kdd/XiaK21, DBLP:conf/iclr/HladisLFT23]. Despite the power of graph models, the relation information in a 3-dimensional space captures the multi-level structure such as the angle between two edges. A line of research works explore the protein structure in 3D space [DBLP:conf/iclr/HermosillaSLFVK21, DBLP:conf/iclr/Huang0ZZZZCW0Y24]. Recently, large language models (LLMs) have also been introduced to model the sequence [DBLP:conf/nips/MeierRVLSR21, lin2023evolutionary] inspired by the success of natural language processing.
Capitalizing on the power of generative models such as Generative Adversarial Networks (GANs) and diffusion models, deep generative modeling has shown its potential for fast generation of new and viable protein structures. [DBLP:conf/nips/AnandH18] has applied GANs to the task of generating protein structures by encoding protein structures in terms of pairwise distances on the protein backbone. Diffusion models [song2021denoising, jiang2023successfully, gao2024image, xiao2025lightcache] have emerged as a powerful tool for graph-structured diffusion processes [DBLP:conf/icml/KlarnerRMDT24]. FrameDiff has been proposed for monomer backbone generation and it can generate designable monomers up to 500 amino acids [DBLP:conf/icml/YimTBMDBJ23]. NOS is another diffusion model that generates protein sequences with high likelihood by taking many alternating steps in the continuous latent space of the model [DBLP:conf/nips/GruverSFRHLRCW23].
@@ Appendix A Appendices
A.1 Experiment settings
Our model leverages the esm2_t33_650M_UR50D model from ESM2 [lin2023evolutionary] as the amino acid-level encoder. To construct the Protein Variational Auto-encoder model, we set the latent dimension to 384, and the decoder is composed of 8 Transformer [vaswani2017attention] decoder blocks, each equipped with an 8-head self-attention mechanism. The Protein Variational Auto-encoder model is trained with a learning rate of $10^{-4}$, using a combination of mean squared error (MSE) and cross-entropy as the loss functions. To regulate the latent vector distribution, we apply a KL divergence loss with a weight of $10^{-5}$. We experimented with 128, 256, and 512 as the maximum sequence lengths. For the diffusion model, we modify the DiT-B architecture from DiT [peebles2023scalable], which consists of 12 DiT blocks and uses a hidden size of 768. The DiT model is trained from scratch with a learning rate of $10^{-4}$ and includes a weight decay of $10^{-5}$.